OpenStudy (anonymous):

a bill in the us house of representatives passed because at least 2/3 of the members present voted in favor of it. if the bill received 284 votes, at least how many members of the house of representives were present for the vote?

OpenStudy (texaschic101):

Is that saying there were 284 votes in favor of the bill ? Or 284 votes total (in favor and against) ?

OpenStudy (texaschic101):

I think it is saying 284 votes in favor of the bill, therefore, there were a total of x voters. 2/3 of what = 284 2/3x = 284 x = 284/(2/3) x = 284 * 3/2 x = 852/2 x = 426. there were 426 voters and 2/3 of them (284 voters) voted for the bill.

OpenStudy (texaschic101):

@Hero .... Am I understanding this problem correctly ?

OpenStudy (texaschic101):

@zepdrix ...take a peek at this please

hero (hero):

Looks good to me.

OpenStudy (texaschic101):

thanks for the confirmation....wasn't sure if it wanted 2/3 of 284 or 2/3 of what= 284

hero (hero):

Well, there would have to be more members present than the number of votes for the bill. If you only found 2/3 of 284, that would not represent the total number of members present.

OpenStudy (texaschic101):

I should have known that...lol....thanks Hero

OpenStudy (anonymous):

yeah